OpenStudy (anonymous):

look i need an answer! Which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring? −b b2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x2 + 7x + 3 = 0

OpenStudy (anonymous):

b^2-4ac if it is a perfect square, then the problem can be solved by factorisation

OpenStudy (anonymous):

so is it 5 ?

OpenStudy (anonymous):

nope. b=7,a=2,c=3 calculate again

OpenStudy (anonymous):

its 25

OpenStudy (richyw):

\[\sqrt{7^2-4(2)(3)}=\sqrt{49-24}=\sqrt{25}\]

OpenStudy (richyw):

so it is 5...

OpenStudy (richyw):

oh sorry

OpenStudy (richyw):

25 is correct.